QUESTION 23 Oroz Company had the following information available: Expected Costs and Selling Price Based on 5,000 units: Variable manufacturing costs per unit $32 Fixed manufacturing costs per unit $20 Selling price per unit $70 Expected production level 5,000 units In the flexible budget at 10,000 units, what is the total manufacturing cost? O $250,000 o $700,000 o $520,000 O $420,000
uppose the supply of product X is perfectly inelastic. If there is an increase in the demand uct X is perfectly inelastic. If there is for this product, equilibrium price A. will decrease, but equilibrium quantity will increase. B. and quantity will both decrease. C. will increase, but equilibrium quantity will decline. will increase, but equilibrium quantity will be unchanged.
Answer #1

Answer : 23) The answer is option C.

Per unit total manufacturing cost = Variable cost + fixed cost = 32 + 20 = $52

For 10,000 units the total manufacturing cost = Per unit total manufacturing cost * Output = 52 * 10,000 = $520,000.

Therefore, option C is correct.

35) The answer is option D.

In case of perfectly inelastic supply if demand for the good increase then the demand curve shift to rightward. As a result, the equilibrium price level increase but the equilibrium quantity level does not change. Therefore, option D is correct.